[Brahms-l] Shift report 20050314 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Mar 14 2005 - 00:06:28 EST
Shifter leader was SB
Shifters were 


1. Beam from previus shift for about 20 minutes. No beam since then for
the whole shift.
2. From minimum bias run from previous shift
FS @6 deg @A861,  MRS @90 deg @B700
run13771  trig4(128k) trig5(474k)

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Mar 14 00:06:47 2005

This archive was generated by hypermail 2.1.8 : Mon Mar 14 2005 - 00:07:02 EST